# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v1

@@ 8 Collaboration between AVs for Prediction
Collaborative perception enhances the capabilities of individual vehicle sensors by utilizing multiple data sources from various angles. By aggregating diverse sensor inputs, it allows for more accurate detection of distant and obscured objects. Depending on the collaboration mode, the existing SOTA models can be grouped into early, intermediate, and late collaboration. Another line of categorization is based on the collaboration task, e.g., prediction [596], planning [597, 598], task-invariant [599]. A substantial part of models in the literature aims to find a balance between increasing the prediction performance and limited bandwidth along with mitigating the impact of communication delays, which can lead to misalignment of the features. Under these challenges, we discuss in this section models, which address transmission delays, communication redundancy and pose misalignment. Lastly, we review the models aiming to collaborate through a unified fusion space addressing the variations in utilized sensors suits and different architectures of prediction modules. This section reviews collaborative models aiming to improve performance not only for prediction but rather various downstream tasks, e.g., detection, segmentation and tracking. Due to the novelty of the field in terms of application to trajectory prediction, reviewing only collaborative prediction models will not provide the reader with the proper overview of the field and available effective fusion mechanism that can be utilized or adapted for collaborative prediction. For more details and challenges on collaborative perception, we refer the reader to comprehensive survey papers [600, 601, 602, 603, 604].
@@ 8.1 Type of Fusion
Collaborative perception can be achieved by combining data and features at different stages of the perception pipeline. Currently, the literature distinguishes three fusion modes: early, intermediate, and late fusion (please refer to Figure 14).
In early fusion, raw data is aggregated at the beginning of the perception pipeline. This approach allows the ego-vehicle to address challenges like occlusions and distance limitations. However, transferring raw data requires significant bandwidth and necessitates efficient suppression techniques to transmit only complementary data, reducing message sizes. Chen et al. [605] proposed an unsupervised 3D representation learning approach (CO3), leveraging point clouds from vehicle and infrastructure side views to generate contrasting yet semantically aligned representations.
In intermediate collaboration, deep semantic features are exchanged between agents, and the ego-vehicle integrates these features to generate final predictions. Compared to early fusion, intermediate fusion involves more information loss and requires optimal feature selection and fusion strategies. Xu et al. [606] proposed V2X-ViT, a vehicle-to-infrastructure cooperative perception system based on a Vision Transformer model. This model employs heterogeneous multi-agent self-attention and multi-scale window self-attention to integrate data, addressing asynchronous data exchange and positioning inaccuracies. Zhou et al. [607] introduced a GNN-based framework for multi-robot perception, evaluated under challenging conditions like camera occlusion and heavy noise. Xu et al. [608] developed CoBEVT, a simple yet high-performing framework for collaborative BEV map creation. Its fused axial attention module (FAX) integrates camera features from various viewpoints and agents, capturing both dense local and broad spatial relationships. Wei et al. [609] introduced CoBEVFlow, a robust framework for collaborative perception using a BEV flow approach to manage non-uniform collaboration messages and transmit perceptual data without generating new features. Yang et al. [610] proposed SCOPE, a framework designed to unify spatio-temporal awareness across agents for effective information exchange. Wang et al. Wang et al. [611] introduced CORE, a bandwidth-efficient framework for multi-agent cooperative perception that combines a compression mechanism, attention-based integration of messages, and a reconstruction module. Lastly, Qiao and Zulkernine [612] developed an adaptive spatial-wise feature fusion model with a trainable feature selection module for efficient information aggregation.
@@
In late fusion, predictions are generated independently by agents and shared at the end of the pipeline. The ego-vehicle combines these outputs to refine its predictions. While this mode is bandwidth-efficient, it can be affected by noise and incomplete data in the individual outputs. Lu et al. [613] proposed a hybrid collaboration framework (CoAlign) to mitigate pose estimation inaccuracies, using an agent-object pose graph to improve alignment and integrating features at multiple spatial resolutions. Yu et al. [614] introduced the DAIR-V2X dataset, a multi-modal and multi-view dataset for vehicle-to-infrastructure collaborative perception. It includes VIC3D, a benchmark model for jointly detecting and classifying 3D objects using vehicular and infrastructural sensors, addressing synchronization issues and transfer costs. Glaser et al. [615] tackled bandwidth limitations and obstructions by introducing the Multi-Agent Spatial Handshaking network, which integrates varying and overlapping raw sensor data from different viewpoints under strict bandwidth constraints.
8.2 Challenges
Current SOTA models not only focus on designing efficient fusion layers but also tackle critical challenges in collaborative perception, such as transmission delays, communication redundancy, and pose misalignment. Communication redundancy manifests in two ways: determining which surrounding agents should share data to complement the ego-vehicle’s observations and designing efficient compression mechanisms to minimize message sizes while maintaining essential information, thereby addressing bandwidth constraints. Transmission latency is another significant challenge, as limited communication speeds and large message sizes can lead to delays and asynchronous data fusion across different timestamps. This issue further complicates data alignment, which is already affected by inaccuracies in localization systems and transformation processes. Effective collaborative frameworks must incorporate mechanisms to handle these delays and misalignments. Lastly, pose misalignment, whether caused by transmission delays or inconsistencies in feature maps, remains a persistent issue. The following sub-sections categorize SOTA approaches based on the specific challenges they address.
Communication Redundancy. Liu et al. [616] proposed a three-stage handshake communication protocol for collaborative perception, addressing the challenge of determining which agents to communicate with. In the first stage, the ego-agent broadcasts a request message to nearby agents. In response, other agents calculate matching scores based on the received request and their own observations. The ego-agent then connects with the agents offering the highest matching scores to obtain the requested feature maps. Several works have focused on addressing bandwidth constraints by designing architectures that fuse the sparsest yet most informative and complementary features [617, 618, 619]. Li et al. [620] introduced a distilled collaboration graph (DiscoGraph), leveraging a teacher-student training methodology. The teacher model integrates early collaboration with comprehensive inputs, while the student model focuses on partial inputs at intermediate stages of collaboration. Hu et al. [621] introduced a spatial confidence map to address the uneven spatial distribution of perceptual data, enabling agents to share only sparse, critical information to enhance complementary perception. Lastly, Chen et al. [622] proposed the TransIFF framework, incorporating a Cross-Domain Adaptation module to align features and a Feature Magnet (FM) module to dynamically fuse instance-level features.
Transmission Delays. Lei et al. [623] highlight the practical challenges of latency in communication systems, which can disrupt collaborative perception. To mitigate these issues, the authors proposed a latency-aware collaborative perception framework. This framework includes SyncNet, a synchronization component that aligns asynchronous sensory data from multiple agents at the feature level. SyncNet employs attention mechanisms and temporal adjustment strategies to synchronize data into a unified timeframe despite delays. Yu et al. [624] introduced the Feature Flow Net (FFNet), which leverages temporal consistency in consecutive frames from infrastructure cameras. FFNet transmits static feature maps through feature-flow fusion, ensuring robust communication and minimizing the impact of latency on collaborative perception.
Pose Misalignment. Vadivelu et al. [625] proposed a model based on V2VNet [596] to address pose noise and its impact on the reliability of spatially dependent communications. To mitigate localization errors, the authors introduced a neural reasoning model capable of facilitating communication, identifying inaccuracies, and collaboratively resolving pose inconsistencies. Gu et al. [626] addressed the challenge of noisy poses and observation misalignments among multiple agents with FeaCo, a system designed to achieve robust feature-level consensus. FeaCo incorporates a Pose-error Rectification Module (PRM) to align feature maps from multiple vehicles and a multi-scale Cross-level Attention Module (CAM) to enhance information aggregation across agents.
@@ 8.2 Challenges
Unified Fusion Space. The literature explores challenges arising from variations in sensor suites, model architectures, and the integration of collaborative components within existing systems. Xu et al. [627] addressed discrepancies in neural network architectures across collaborating agents by proposing a trainable feature resizer to harmonize feature dimensions and a sparse cross-domain transformer for domain adaptation. In the related study [628], the authors introduced a confidence calibrator to adjust for biases in prediction confidence scores across agents. Xiang et al. [629] tackled the challenge of heterogeneous sensor modalities with a model for hetero-modal cooperative perception, enabling collaboration among agents using different sensor suites. By removing the assumption of uniform sensor setups, the model supports new, diverse agent types in collaborative perception. Lu et al. [630] proposed the HEAL framework, which constructs a unified feature space using a pyramid fusion network and aligns new agent types to this space via backward alignment. Li et al. [631] addressed discrepancies in LiDAR data with a model comprising three components: a Domain-Mixing Instance Augmentation module to ensure uniform data representation, a Progressive Domain-Invariant Distillation module for domain-neutral feature adaptation, and a Domain-Adaptive Fusion module with calibration-aware attention mechanisms. Finally, Ma et al. [632] introduced the MACP framework, equipping single-agent models with cooperative capabilities by freezing the existing prediction model and integrating trainable lightweight modules.
@@ 8.3 Benchmark Datasets
8.4 SOTA Prediction Models
For benchmarking and evaluation of new collaborative prediction models, Table 22 summarizes existing collaborative prediction models detailing the type of fusion and dataset on which the model was evaluated.
V2VNet [596] is a network designed for joint perception and prediction. Its architecture comprises three components: first, a CNN block processes raw sensor inputs to generate a compressible intermediate form; next, a cross-vehicle integration mechanism combines the received data with the ego-vehicle’s data; and finally, the fused intermediate output is used to generate the final predictions. Robust V2VNet [625], an extension of V2VNet, focuses on addressing localization errors for improved performance. V2X-Graph [640] is a late fusion framework that facilitates collaborative trajectory forecasting using graph-based feature fusion. The framework constructs a graph where agents’ motion features and map elements are represented as nodes, and their spatial-temporal relationships as edges. It employs three interpretable sub-graphs: the Agent-Lane sub-graph (ALG) for map-based interactions, the Motion Fusion sub-graph (MFG) for motion features, and the Cooperative Interaction sub-graph (CIG) for inter-agent relationships. These sub-graphs aggregate heterogeneous features across views to enhance trajectory prediction accuracy. The V2XFormer model [638] is a V2X-based framework that utilizes a Swin Transformer as its backbone to extract BEV (Bird’s Eye View) features from multi-view camera data. These BEV features, generated by intelligent vehicles and infrastructure, are spatially aligned to the ego-vehicle’s coordinate system and fused through CoBEVT [608] to produce aggregated representations. The fused BEV features are subsequently used for tasks such as motion prediction, 3D object detection, and accident prediction. Lastly, the CMP framework [641] enhances trajectory prediction by fusing data at multiple stages. During perception, BEV features are compressed and shared for fusion using the FuseBEVT module [608], improving object detection and tracking. In the prediction aggregation stage, vehicles share their predicted trajectories, which are further refined through a multi-head attention mechanism for enhanced trajectory prediction accuracy.
@@ 8.5 Summary and Discussion
Collaborative perception overcomes limitations in individual perception systems, such as restricted sensing ranges, occlusions in dense environments, and sensor noise. However, its implementation poses significant challenges for collaborative trajectory prediction, including pose misalignment, communication redundancy, and time delays, which remain open issues. This section examines different fusion approaches, reviews SOTA models addressing these challenges, and summarizes collaborative datasets alongside descriptions of trajectory prediction SOTA models.
The concept of multiple vehicles sharing complementary information assumes a perfectly functioning, fully connected collaboration network. Most existing models operate under the assumption that the ego-vehicle aggregates information while all other nodes act as transmitters. A key research direction is exploring diverse communication topologies, transmitting frequencies and naturally occurring time delays within such networks. Current models primarily address bandwidth limitations but often overlook the impact of transmission delays. In dynamic tasks like trajectory prediction, data transmitted even 0.5 seconds late may become irrelevant for the ego-vehicle, underscoring the critical need to account for these delays. Another research direction addresses the assumption in existing models that the ego-vehicle performs fusion and decodes trajectories. However, none of the current studies provide an inference time analysis. In trajectory prediction, timely delivery is critical, as planning must receive predictions within a specific time window. Providing fused features could significantly slow down inference. Beyond time complexity analysis, it is essential to develop a smart strategy to determine when fusion is necessary and when to rely solely on existing individual predictions. In a collaborative mode, the assumption is that the system operates as a network of intelligent agents, each capable of autonomous operation and fusing information only when required.
Lastly, in a network of intelligent vehicles, the reliability and trustworthiness of shared information must be thoroughly evaluated before fusion. Autonomous vehicles perceive their environment using both on-board sensors and data received from other vehicles within the network. However, the accuracy of this perception depends not only on the performance of an individual vehicle’s sensors but also on the quality of the shared data. A transmitting vehicle may have malfunctioning or degraded sensors, introducing errors into the cooperative system. To address this, a verification step should be implemented to assess the credibility of incoming data, ensuring that only trustworthy information contributes to the fused representation. This process is critical to maintaining the integrity and effectiveness of the collaborative perception system.
@@ 9 Discussion
In recent years, both academic and industrial research communities have made significant progress in advancing trajectory prediction methods, characterized by diverse approaches and innovative solutions. This section summarizes the topics reviewed and addresses three key questions posed at the outset: (Q1) What are the existing solutions for trajectory prediction? Section 9.1 revisits available modelling approaches in the literature to answer this question. (Q2) What are the emerging research trends in trajectory forecasting realm? Section 9.2 discusses the primary focus areas of researchers, the challenges being tackled, and emerging trends. (Q3) Is the trajectory prediction problem solved, or is there still room for improvement? Section 9.3 talks about existing challenges and gaps in current studies. Finally, Section 9.4 elaborates more on open challenges and proposes potential future research directions.

### v2

@@ 8 Collaboration between AVs for Prediction
There are certain limitations that cannot be resolved through multimodal inputs or advanced mechanisms. Each sensor has inherent constraints, e.g., it is limited in range, can get miscalibrated. Moreover, even for human drivers, occlusions are a common cause of accidents; for autonomous vehicles (AVs), they present an even greater challenge, as AVs lack human intuition and their perception of the environment relies entirely on sensors. When the limitations of an individual sensor suite are reached, collaboration between vehicles can compensate by aggregating information from multiple viewpoints. The current SOTA work in this area is primarily focused on collaborative perception, where sensor data from multiple vehicles is fused to construct a holistic view of the environment. This fused representation then supports downstream tasks such as object detection [593, 594], tracking [595], trajectory prediction [596], planning [597, 598], and task-invariant modules [599].
A major research focus has been the design of efficient fusion architectures to enhance the accuracy of a down-stream task. Based on the stage at which information is integrated, existing models are commonly classified into early, intermediate, and late fusion. Each mode imposes different bandwidth requirements and communication overhead. Communication redundancy has thus become a central research direction, with methods ranging from sparse intermediate fusion to selective sharing of complementary information. These strategies aim to reduce bandwidth consumption while preserving critical context that complements ego-vehicle observations. Other lines of work address bandwidth optimization [600, 594, 601, 602], mitigation of communication delays [603, 604], and handling pose misalignment [605, 606, 607]. Additionally, due to different sensor configurations across vehicles, recent models aim to learn in a unified fusion space [608, 609, 610, 611], addressing both heterogeneous sensor suites and varying perception architectures.
This section reviews existing collaborative models, surveys available multi-vehicle datasets, and discusses current limitations along with future research directions. It is important to note that the categorization here is based on fusion strategies, rather than the collaboration type (e.g., vehicle-to-vehicle or vehicle-to-infrastructure). For a broader overview of inter-vehicle collaboration challenges, we refer readers to recent comprehensive surveys [612, 613, 614, 615, 616].
@@ 8.1 Type of Fusion
Collaborative between vehicles can be achieved by combining data and features at different stages. Currently, the literature distinguishes three fusion modes: early, intermediate, and late fusion (please refer to Figure 14).
In early fusion, raw data is combined at the beginning of the perception pipeline, allowing the ego vehicle to utilize complementary sensor information to mitigate occlusions. Chen et al. [593] introduced an unsupervised 3D representation learning approach, CO3, which uses point clouds from both vehicle and infrastructure side views to generate contrasting and semantically aligned representations. Since transmitting raw data demands considerable bandwidth, Li et al. [594] introduced a distilled collaboration graph (DiscoGraph), leveraging a teacher-student training methodology. The teacher model integrates early collaboration with comprehensive inputs, while the student model focuses on partial inputs at intermediate stages of collaboration. Building on the combination of early and intermediate fusion, Li et al. [617] investigated domain generalization in collaborative perception and introduced Cooperative Mixup Augmentation for Generalization (CMAG) to enhance model generalization by emulating unseen collaborative scenarios.
In intermediate fusion [618, 619], deep semantic features are shared, which are further integrated by the ego-vehicle for the downstream task. Compared to early fusion, intermediate fusion involves more information loss and requires optimal feature selection and fusion strategies. Xu et al. [620] proposed V2X-ViT, a vehicle-to-infrastructure cooperative perception system based on a Vision Transformer model. This model employs heterogeneous multi-agent self-attention and multi-scale window self-attention to integrate data, addressing asynchronous data exchange and positioning inaccuracies. Zhou et al. [595] introduced a GNN-based framework for multi-robot perception, evaluated under challenging conditions like camera occlusion and heavy noise. Xu et al. [621] developed CoBEVT, a simple yet high-performing framework for collaborative BEV map creation. Its fused axial attention module (FAX) integrates camera features from various viewpoints and agents, capturing both dense local and broad spatial relationships. Wei et al. [622] introduced CoBEVFlow, a robust framework for collaborative perception using a BEV flow approach to manage non-uniform collaboration messages and transmit perceptual data without generating new features. Yang et al. [623] proposed SCOPE, a framework designed to unify spatio-temporal awareness across agents for effective information exchange. Wang et al. Wang et al. [624] introduced CORE, a bandwidth-efficient framework for multi-agent cooperative perception that combines a compression mechanism, attention-based integration of messages, and a reconstruction module. Lastly, Qiao and Zulkernine [625] developed an adaptive spatial-wise feature fusion model with a trainable feature selection module for efficient information aggregation.
@@
In late fusion, the decisions are generated independently by each vehicle and further shared. The ego-vehicle aggregates these outputs to refine its own predictions. While this mode is bandwidth-efficient, it is susceptible to noise and incomplete information in the individual predictions. CoAlign [626] proposes a hybrid collaboration framework that combines intermediate and late fusion, mitigating pose estimation inaccuracies by constructing an agent-object pose graph for improved alignment and integrating features at multiple spatial resolutions. Yu et al. [627] introduced the DAIR-V2X dataset along with the benchmark late fusion model VIC3D, which compensates for temporal asynchrony by estimating object motion from consecutive frames and projecting states to the vehicle’s current time before fusing the adjusted predicted boxes.
8.2 Communication Redundancy
Mitigating communication redundancy has been explored in two main directions: (1) determining which information should be shared by surrounding agents to complement the ego vehicle’s observations [600, 628, 629, 630, 631], and (2) designing efficient compression mechanisms to minimize message sizes while preserving essential information [601, 602]. The first line of research typically follows a standard pipeline: the ego agent generates and broadcasts a request message, surrounding agents evaluate matching scores or generate feature maps based on both their local observations and the received request, and then transmit the selected responses. The resulting messages, when aggregated by the ego vehicle, offer complementary information that enhances perception. In the second category, Hu et al. [601] introduced a spatial confidence map that enables surrounding agents to transmit only sparse, high-confidence features, thus improving efficiency while maintaining perception quality. Chen et al. [602] proposed TransIFF, a framework that filters out low-confidence predictions to reduce message size and includes a Cross-Domain Adaptation module for aligning features, as well as a Feature Magnet (FM) module to dynamically fuse instance-level features. Hu et al. [632] presented a two-stage communication reduction method: first, applying a codebook-based message representation for compression, and second, using an information-filling strategy to prioritize the most informative content for transmission. Along the same lines, Ding et al. [633] proposed a compact object-level representation that preserves both semantic and structural information through a learnable sampling strategy.
8.3 Transmission Delays
Transmission latency is another challenge that arises when vehicles collaborate. Even under the assumption that all collaborators share the same parameters, such as FPS and broadcasting frequencies, delays are inevitable, since no information share happens instantly. The delay issue further complicates data alignment, which is already affected by inaccuracies in localization systems and transformation processes. Lei et al. [603] proposed a latency-aware collaborative perception framework with a synchronization component called SyncNet, which uses attention mechanisms and temporal adjustment strategies to synchronize data into a unified timeframe. Yu et al. [604] introduced the Feature Flow Net (FFNet), which leverages temporal consistency from consecutive frames captured by infrastructure cameras to derive a feature flow. This flow serves as a predictive function to describe feature changes over time, generating features aligned with the current timeframe of ego-vehicles. Xu et al. [634] proposed a cooperative framework in which BEV features are extracted with uncertainty quantification and aligned by warping them using motion displacement to match the ego-agent’s timestamp.
8.4 Unified Fusion Space
@@
The literature explores challenges arising from variations in sensor suites, model architectures, and the integration of collaborative components within existing systems. One category of studies proposes collaboration through a unified fusion space [608, 609, 610, 611], addressing variations in sensor suites and prediction module architectures. Xu et al. [635] addressed discrepancies in neural network architectures across collaborating agents by proposing a trainable feature resizer to harmonize feature dimensions and a sparse cross-domain transformer for domain adaptation. In the related study [636], the authors introduced a confidence calibrator to adjust for biases in prediction confidence scores across agents. Xiang et al. [608] tackled the challenge of heterogeneous sensor modalities with a model for hetero-modal cooperative perception, enabling collaboration among agents using different sensor suites. By removing the assumption of uniform sensor setups, the model supports new, diverse agent types in collaborative perception. Lu et al. [637] proposed the HEAL framework, which constructs a unified feature space using a pyramid fusion network and aligns new agent types to this space via backward alignment. Li et al. [609] addressed discrepancies in LiDAR data with a model comprising three components: a Domain-Mixing Instance Augmentation module to ensure uniform data representation, a Progressive Domain-Invariant Distillation module for domain-neutral feature adaptation, and a Domain-Adaptive Fusion module with calibration-aware attention mechanisms. Finally, Ma et al. [610] introduced the MACP framework, equipping single-agent models with cooperative capabilities by freezing the existing prediction model and integrating trainable lightweight modules.
@@ 8.5 Benchmark Datasets
8.6 Collaborative Trajectory Prediction Models
Table 22 summarizes SOTA collaborative prediction models detailing the type of fusion and dataset on which the models were evaluated. V2VNet [596] employs intermediate fusion, using a CNN block to process raw sensor inputs into a compressible intermediate representation. A cross-vehicle integration mechanism then fuses these intermediate outputs for collaborative prediction. In BEV-centered fusion, V2XFormer [643] utilizes a Swin Transformer to generate BEV features, which are spatially aligned to the ego-vehicle’s coordinate system. These aligned features are further fused using CoBEVT [621] to produce aggregated representations that are subsequently decoded into motion predictions. Leveraging graph neural networks (GNNs), V2X-Graph [645] constructs a graph from observed trajectories and vector maps, representing agents’ motion features and map elements as nodes, with spatial-temporal relationships as edges. It employs three interpretable sub-graphs: the Agent-Lane sub-graph (ALG) for map-based interactions, the Motion Fusion sub-graph (MFG) for motion features, and the Cooperative Interaction sub-graph (CIG) for inter-agent relationships. These sub-graphs aggregate heterogeneous features across views, which are then decoded into future trajectories. Next, a hybrid fusion prediction model, the CMP model [646], enhances trajectory prediction through two stages of fusion. In the perception stage, it performs intermediate fusion by using the FuseBEVT module [621] to improve object detection. In the prediction stage, it applies late fusion, refining predictions through a multi-head attention mechanism. Lastly, Co-MTP [647] constructs a heterogeneous graph by integrating past and future trajectories from both ego and infrastructure perspectives. It applies a graph transformer for intermediate fusion in the history domain to enhance past features, and leverages future trajectories to model interactions that may influence ego planning.
@@ 8.7 Summary and Discussion
Collaboration between vehicles can help overcome key limitations in individual perception systems, such as restricted sensing range, occlusions in dense traffic, and sensor degradation. This section reviewed different fusion strategies, highlighted SOTA collaborative prediction models, and summarized available multi-vehicle datasets.
Currently, collaboration is mainly explored at the perception level, which poses significant challenges due to high bandwidth requirements. While constructing a unified global view and decoding it downstream is a well-justified approach, shifting fusion to task-specific stages, such as trajectory prediction, may offer more flexibility and compliance with existing communication protocols requirements. Prediction models typically rely on structured, low-dimensional inputs (e.g., past trajectories, HD maps, and interaction features), rather then dense feature maps shared in collaborative perception. By moving fusion to the prediction stage, collaboration can become both semantically more relevant to the trajectory forecasting task and communication-efficient. This direction remains underexplored.
Next, most existing frameworks assume a fully connected network where the ego vehicle receives data from all collaborators. However, practical deployments must account for varying communication topologies, transmission frequencies, and delays. In trajectory prediction, even a half second delay may render shared data obsolete. Thus, addressing time sensitivity and synchronization is critical. Another limitation is the lack of inference time analysis in current models. Prediction must be delivered within tight time windows to be actionable for planning. Transmitting and fusing high-dimensional features may hinder latency requirements. Additionally, verifying the trustworthiness of shared information is crucial for accuracy. A vehicle with degraded sensors can unintentionally introduce errors into the system. Fusion mechanisms must incorporate verification stages to assess the reliability of incoming data and prevent the propagation of low-quality inputs.
Finally, the development of standardized testing platforms is essential for prototyping and benchmarking collaborative algorithms. OpenCDA [648] is one such open-source ecosystem, offering tools for cooperative detection and automation testing. Expanding such platforms to cover V2I collaboration, communication robustness, and scenario diversity will be important for advancing the field.
@@ 9 Discussion
In recent years, both academic and industrial research communities have made significant progress in advancing the field of trajectory prediction. To summarize these advancements, this section revisits the research questions posed at the beginning of the survey, elaborates on the remaining open challenges, and outlines potential future research directions.